OpenStudy (anonymous):

Over two summers, Ray saved $900 and $500. The polynomial 900x^2 + 500x represents his savings at the beginning of the third year, where x is the growth factor. (The interest rate r is x - 1.) What is the interest rate he needs to start the third year with $1600 in savings? A. 1.1% B. 0.1% C. 8.4% D. 264% my answer is C can someone plz tell me if it's correct

OpenStudy (anonymous):

\[ 900x^2 + 500x = 1600 \Rightarrow 900x^2 + 500x - 1600 = 0 \] Use the quadratic equation to solve for \(x\). You will have two solutions. Pick the most logical one (one that isn't negative. Plug that into \(x\) for \(r = x - 1\) to get \(r\).

OpenStudy (anonymous):

ok so its A

OpenStudy (anonymous):

No. C is the correct answer. When we solve for \(x\) we get \(1.0842\) and a negative solution we can throw away. So \(r = 1.0842 - 1 = 0.0842\). Now we change this into a percent and get \(8.42 \% \)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I never meant to imply that C is incorrect. I just wanted to make sure you knew the method.

OpenStudy (anonymous):

oh ok thank you you have been a big help